3 red 3 yellow 3 green and 3 white marbles what is the probability of picking the 3 red before you pick 3 of any other color 10 times in a row then 20 times in a row is there anyine out there that can answer this thank you
Answer by ewatrrr(24785)


Hi
12 marbles: 3R, 3Y, 3G, 3W
P(PICKING A RED) = 3/12 = 1/4
(WITH REPLACEMENT) 10 TIMES IN A ROW
(1/4)^10 | (1/4)(1/4)(1/4)(1/4)(1/4)(1/4)(1/4)(1/4)(1/4)(1/4)
(WITH REPLACEMENT) 20 TIMES IN A ROW
(1/4)^20
Wish You the Best in your Studies.